Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Ring Heater)
lumaca - 14:56 Saturday 03 June 2023 (60433) Print this report
NE RH switched OFF

After the ENV morning activity, at 12.23 UTC NE RH has been switched OFF (Fig. 1), in view of NE tower venting.

The NE RH must be cold (INF_TCS_NE_RH_TE < 20°C) with a tower pressure below 10e-2 mbar before starting the venting, to avoid oxidation issues.

 

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×